# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ 8.4 Proofs in Section 5
Let $K,\Sigma$ fixed and let $f_{K}$ be defined in (16). Observe that $\Sigma^{\prime}$ defined in (25) is the maximizer of $f_{K}$. Then by Lemma 3.3,
@@
where $(a)$ follows the fact that $f_{K}(\Sigma^{\prime})-f_{K}(\Sigma)\geq 0$ and $\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\leq 1$; $(b)$ follows from fact that $f_{K}(\Sigma^{\prime})\geq f_{K}(\Sigma^{*}).$ Finally, apply (48) to get
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)\leq$ $-\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\left(C(K,\Sigma)-C(K^{*},\Sigma^{*})\right).$
@@
Let $K^{\prime}$ denote $K^{(t+1)}$, $K$ denote $K^{(t)}$ and $S^{\prime}$ denote $S^{(t+1)}$. Apply Lemma 3.3 with $f_{K}$ defined in (16) to get
@@
where $(a)$ follows from the assumption $\|S^{\prime}-S^{*}\|\leq\sigma_{\min}(S^{*})$ which implies $-1+\left\|(S^{\prime}-S^{*})S^{*-1}\right\|\in[-1,0]$ and the fact $\Sigma^{\prime}$ is the maximizer of $f_{K}.$ Finally, apply (48) to get
@@
$\mathcal{F}_{K}(X)$ $:=(A-BK)X(A-BK)^{\top},\quad\mathcal{G}^{K}_{t}(X):=(A-BK)^{t}X(A-BK)^{\top t}$ (57)
$\mathcal{T}_{K}(X)$ $:=\sum_{t=0}^{\infty}\gamma^{t}(A-BK)^{t}X(A-BK)^{\top t}=\sum_{t=0}^{\infty}\gamma^{t}\mathcal{G}^{K}_{t}(X).$
@@
$\mathcal{T}_{K}=(I-\gamma\mathcal{F}_{K})^{-1}.$ (58)
We also define the induced norm for these operators as $\|T\|:=\sup_{X}\frac{\|T(X)\|}{\|X\|},$ where $T=\mathcal{F}_{K},\mathcal{G}_{t}^{{K}},\mathcal{T}_{{K}}$ and the supremum is over all symmetric matrix $X\in\mathbb{R}^{n\times n}$ with non-zero spectral norm.
@@
For any $K\in\mathbb{R}^{k\times n}$ and any symmetric $\Sigma\in\mathbb{R}^{n\times n}$,
$S_{K,\Sigma}=\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\mathcal{G}^{K}_{t-s}(B\Sigma B^{\top}+W).$
Proof.
@@
$\mathbb{E}[x_{t}x_{t}^{\top}]=(A-BK)^{t}\mathbb{E}[x_{0}x_{0}^{\top}](A-BK)^{\top t}+\sum_{s=1}^{t}(A-BK)^{t-s}(B\Sigma B^{\top}+W)(A-BK)^{\top t-s}.$
@@
Thus, the induction is completed. Then by (57), we can write $S_{K,\Sigma}$ as:
$S_{K,\Sigma}=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[x_{t}x_{t}^{\top}]=\mathcal{T}_{K}\left(\mathbb{E}[x_{0}x_{0}^{\top}]\right)+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\mathcal{G}_{t-s}^{K}(B\Sigma B^{\top}+W).$
@@
For any $K_{1}$ and $K_{2}$,
$\left\|\mathcal{F}_{K_{1}}-\mathcal{F}_{K_{2}}\right\|\leq(\|A-BK_{1}\|+\|A-BK_{2}\|)\|B\|\left\|K_{1}-K_{2}\right\|.$
Lemma 8.4 follows a direct calculation with triangle inequality.
@@
For any $K_{1}$ and $K_{2}$ satisfying $\|A-BK_{1}\|\leq\rho$ and $\|A-BK_{2}\|\leq\rho$, with $\xi_{\gamma,\rho}$ defined in (26),
$\left\|\left(\mathcal{T}_{K_{1}}-\mathcal{T}_{K_{2}}\right)(X)\right\|$ $\leq\sum_{t=0}^{\infty}\gamma^{t}\left\|\left(\mathcal{G}_{t}^{K_{1}}-\mathcal{G}^{K_{2}}_{t}\right)(X)\right\|\leq\xi_{\gamma,\rho}\|\mathcal{F}_{K_{1}}-\mathcal{F}_{K_{2}}\|\|X\|,$ (59)
$\sum_{t=0}^{T-1}\|(\mathcal{G}_{t}^{\prime}-\mathcal{G}_{t})(X)\|$ $\leq\frac{2-\rho^{2}-\rho^{2T}}{(1-\rho^{2})^{2}}\|\mathcal{F^{\prime}}-\mathcal{F}\|\|X\|,\quad\forall\,T\geq 1.$ (60)
Proof.
@@
$\left\|\left(\mathcal{G}_{t+1}^{\prime}-\mathcal{G}_{t+1}\right)(X)\right\|$ $=\left\|\mathcal{F}^{\prime}\circ\mathcal{G}_{t}^{\prime}(\Sigma)-\mathcal{F}^{\prime}\circ\mathcal{G}_{t}(X)+\mathcal{F}^{\prime}\circ\mathcal{G}_{t}(X)-\mathcal{F}\circ\mathcal{G}_{t}(X)\right\|$ (61)
@@
Summing (61) for $t=0,1,2,\cdots$ with a discount factor $\gamma$, we have
@@
The proof of (59) is finished by noting $\mathcal{G}_{0}=\mathcal{F}$ and $\mathcal{G}^{\prime}_{0}=\mathcal{F}^{\prime}$.
To show (60), fix integer $T\geq 1.$ Summing up (61) for $t=0,1,\cdots,T-1$ to get
$\sum_{t=0}^{T-1}\left\|\left(\mathcal{G}_{t+1}^{\prime}-\mathcal{G}_{t+1}\right)(X)\right\|\leq\rho^{2}\sum_{t=0}^{T-1}\left\|\left(\mathcal{G}_{t}^{\prime}-\mathcal{G}_{t}\right)(X)\right\|+\frac{1-\rho^{2T}}{1-\rho^{2}}\left\|\mathcal{F}^{\prime}-\mathcal{F}\right\|\|X\|.$ (62)
Note that left-hand side equals to
@@
Plugging the above equation into (62) finishes the proof of (60). ∎
Proof.
(of Lemma 5.5). Denote $\mathcal{G}_{t}=\mathcal{G}^{{K_{1}}}_{t},\mathcal{G}^{\prime}_{t}=\mathcal{G}^{{K_{2}}}_{t},\mathcal{T}=\mathcal{T}_{K_{1}},\mathcal{T}^{\prime}=\mathcal{T}_{K_{2}},\mathcal{F}=\mathcal{F}_{K_{1}}$ and $\mathcal{F}^{\prime}=\mathcal{F}_{K_{2}}$ to ease the exposition. First observe that for any $X,X^{\prime}\in\mathbb{R}^{n\times n}$ and $t\geq 0,$
$\|\mathcal{G}_{t}(X)-\mathcal{G}_{t}(X^{\prime})\|$ $=\|(A-BK)^{t}(X-X^{\prime})(A-BK)^{\top t}\|\leq\rho^{2t}\|X-X^{\prime}\|.$ (63)
By Lemma 8.3,
$\|S_{K_{1},\Sigma_{1}}-S_{K_{2},\Sigma_{2}}\|\leq\|(\mathcal{T}-\mathcal{T}^{\prime})(\mathbb{E}[x_{0}x_{0}^{\top}])\|+\|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}(\mathcal{G}_{t-s}^{{}^{\prime}}-\mathcal{G}_{t-s})(B\Sigma_{1}B^{\top}+W)\|$
$\quad\quad+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\|\mathcal{G}_{t-s}^{\prime}(B\Sigma_{1}B^{\top}+W)-\mathcal{G}_{t-s}^{\prime}(B\Sigma_{2}B^{\top}+W)\|$
$\overset{(a)}{\leq}\|(\mathcal{T}-\mathcal{T}^{\prime})(\mathbb{E}[x_{0}x_{0}^{\top}])\|+\|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=0}^{t-1}(\mathcal{G}_{s}^{{}^{\prime}}-\mathcal{G}_{s})(B\Sigma_{1}B^{\top}+W)\|$
$\quad\quad+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\rho^{2(t-s)}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|$
$\overset{(b)}{\leq}\xi_{\gamma,\rho}\|\mathcal{F}-\mathcal{F}^{\prime}\|\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\sum_{t=0}^{\infty}\gamma^{t}\frac{2-\rho^{2}-\rho^{2t}}{(1-\rho^{2})^{2}}\|\mathcal{F^{\prime}}-\mathcal{F}\|\|B\Sigma_{1}B^{\top}+W\|+\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|$
$=\xi_{\gamma,\rho}\|\mathcal{F}-\mathcal{F}{{}^{\prime}}\|\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\|\mathcal{F^{\prime}}-\mathcal{F}\|\cdot\|B\Sigma_{1}B^{\top}+W\|+\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|,$
where $(a)$ follows from (63); $(b)$ follows from (60); and Lemma 8.5. The constants $\xi_{\gamma,\rho},\zeta_{\gamma,\rho},\omega_{\gamma,\rho}$ are defined in (26). Finally, applying Lemma 8.4 finishes the proof. ∎
@@
The objective of this section is to provide a bound for $\|S^{(t+1)}-S^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$, as summarized Lemma 5.6:
Note that Lemma 5.5 in Appendix 8.4.3 can be employed to derive a bound on $\|S^{(t+1)}-S^{*}\|$ in relation to $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|.$ In this section, we further establish this bound by deriving the bounds for $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$ and $\|P_{K^{(t)}}-P_{K^{*}}\|$ (cf. Lemma 8.6 and Lemma 8.7). Additionally, the perturbation analysis for $P_{K}$ in Lemma 8.8 demonstrates $\|P_{K^{(t)}}-P_{K^{*}}\|$ can be bounded by $\|{K^{(t)}}-{K^{*}}\|$, which completes the proof for Lemma 5.6.
We first establish the bound for the one-step update of $K$ and $\Sigma$ in Lemma 8.6 and Lemma 8.7 respectively.
@@
$\|K^{\prime}-K^{*}\|$ $=$ $\|K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}-K^{*}+(R+\gamma B^{\top}P_{K}B)^{-1}E_{K^{*}}\|$ (65)
@@
$\|E_{K}-E_{K^{*}}\|$$\leq$ $\gamma\|B\|\|A\|\|P_{K}-P_{K^{*}}\|+\|(\gamma B^{\top}P_{K^{*}}B+R)K^{*}-(\gamma B^{\top}P_{K^{*}}B+R)K\|$ (66)
@@
Combining (65) and (66), then using $\kappa\geq\|K\|$ for any $K\in\Omega$ completes the proof. ∎
@@
Proof.
(of Lemma 8.7). Observe from (IPO) and (6) that
@@
Lemma 8.8 perform perturbation analysis on $P_{K}$ and establish bounds for the difference in $P_{K}$ with respect to the perturbation in $K$. Consequently, both $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ (in Lemma 8.6 and Lemma 8.7) can be bounded in terms of $\|K^{(t)}-K^{*}\|$.
@@
For any $K\in\Omega,$ with $c$ defined in (27), $\|P_{K}-P_{K^{*}}\|\leq c\|K-K^{*}\|.$
Proof.
(of Lemma 8.8). Fix $K\in\Omega.$ By (58) and (17),
@@
$\|P_{K}-P_{K^{*}}\|\leq\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|+\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|.$ (67)
For the first term in (67), we can apply Lemma 8.5 and Lemma 8.4 to get
$\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|$ (68)
@@
For the second term in (67), note that by Lemma 17 in [12], $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\|\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\|.$ Since $S_{K,\Sigma}\succeq\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])$, thus $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\sigma_{max}\left(\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\right)\leq\frac{1}{\mu}\|S_{K,\Sigma}\|.$ Then we have
$\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|$ (69)
@@
Plugging (68), (69), and (27) in (67) completes the proof. ∎
@@
Proof.
(of Lemma 5.6). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$ to ease the notation. Apply Lemma 8.8 with $K$ and $K^{*}$ to get
$\|P_{K^{*}}-P_{K}\|\leq c\left\|K-K^{*}\right\|.$ (70)
With the assumption that $\|A-BK^{\prime}\|\leq\rho$, $\|A-BK^{*}\|\leq\rho$, we can apply Lemma 5.5 to obtain:
$\|S_{K^{*},\Sigma^{*}}-S_{K^{\prime},\Sigma^{\prime}}\|\leq$ $\left(\xi_{\gamma,\rho}\cdot\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\cdot\|B\Sigma^{*}B^{\top}+W\|\right)\cdot 2\rho\|B\|\left\|K^{*}-K^{\prime}\right\|$ (71)
$\quad+\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma^{*}-\Sigma^{\prime}\|.$
Apply Lemma 8.6 with $\|K\|\leq\kappa$ and (70) to get
$\|K^{\prime}-K^{*}\|$
$\leq$ $\left(1+\sigma_{\min}(R)\|\gamma B^{\top}P_{K^{*}}B+R\|\right)\|K-K^{*}\|+\gamma\sigma_{\min}(R)\left(\|B\|\|A\|+\|B\|^{2}\kappa\right)\|P_{K}-P_{K^{*}}\|$
[…]
Apply Lemma 8.7 with (70) to get
$\|\Sigma^{\prime}-\Sigma^{*}\|\leq\frac{\tau\gamma\|B\|^{2}}{2\sigma_{\min}(R)^{2}}\|P_{K}-P_{K^{*}}\|\leq c\frac{\tau\gamma\|B\|^{2}}{2\sigma_{\min}(R)^{2}}\|{K}-{K^{*}}\|.$
Finally, plugging the above two inequalities into (71) finishes the proof. ∎

### v4

@@ 9.5 Proofs in Section 5
Fix $K,\Sigma$ and let $f_{K}$ be defined in (3.4). Observe that $\Sigma^{\prime}$ following (IPO) update is the maximizer of $f_{K}$. Then by Lemma 3.3,
@@
where $(a)$ follows the fact that $f_{K}(\Sigma^{\prime})-f_{K}(\Sigma)\geq 0$ and $0<\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\leq 1$; $(b)$ follows from fact that $f_{K}(\Sigma^{\prime})\geq f_{K}(\Sigma^{*}).$ Finally, apply (9.11) to get $C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)\leq-\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\left(C(K,\Sigma)-C(K^{*},\Sigma^{*})\right)$ and add $C(K,\Sigma)-C(K^{*},\Sigma^{*})$ to both sides finishes the proof.
@@
Let $K^{\prime}$ denote $^{(t+1)}, KK$ denote $K^{(t)}$ and $S^{\prime}$ denote $S^{(t+1)}$. Apply Lemma 3.3 with $f_{K}$ defined in (3.4) to get
@@
where $(a)$ follows from the assumption $\|S^{\prime}-S^{*}\|\leq\sigma_{\min}(S^{*})$ which implies $1+\left\|(S^{\prime}-S^{*})S^{*-1}\right\|\in[-1,0]$ and the fact $\Sigma^{\prime}$ is the maximizer of $f_{K}.$ Finally, by (9.11),
@@
(9.21) $\mathcal{F}_{K}(X)$ $=(A-BK)X(A-BK)^{\top},\quad\mathcal{G}^{K}_{t}(X)=(A-BK)^{t}X(A-BK)^{\top t}$
$\mathcal{T}_{K}(X)$ $=\sum_{t=0}^{\infty}\gamma^{t}(A-BK)^{t}X(A-BK)^{\top t}=\sum_{t=0}^{\infty}\gamma^{t}\mathcal{G}^{K}_{t}(X).$
@@
(9.22) $\mathcal{T}_{K}=(I-\gamma\mathcal{F}_{K})^{-1}.$
We also define the induced norm for these operators as $\|T\|:=\sup_{X}\frac{\|T(X)\|}{\|X\|},$ where $T=\mathcal{F}_{K},\mathcal{G}_{t}^{{K}},\mathcal{T}_{{K}}$ and the supremum is over all symmetric matrix $X\in\mathbb{R}^{n\times n}$ with non-zero spectral norm. We also define the induced norm for these operators as $\|T\|:=\sup_{X}\frac{\|T(X)\|}{\|X\|},$ where $T=\mathcal{F}_{K},\mathcal{G}_{t}^{{K}},\mathcal{T}_{{K}}$ and the supremum is over all symmetric matrix $X\in\mathbb{R}^{n\times n}$ with non-zero spectral norm.
@@
$S_{K,\Sigma}=\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\mathcal{G}^{K}_{t-s}(B\Sigma B^{\top}+W),$ for any $K,\Sigma\in\Omega.$
Proof 9.8.
@@
$\mathbb{E}[x_{t}x_{t}^{\top}]$ $=(A-BK)^{t}\mathbb{E}[x_{0}x_{0}^{\top}](A-BK)^{\top t}$
$\quad+\sum_{s=1}^{t}(A-BK)^{t-s}(B\Sigma B^{\top}+W)(A-BK)^{\top t-s}.$
@@
Thus, the induction is completed. The proof is finished by noting (9.21).
@@
$\left\|\mathcal{F}_{K_{1}}-\mathcal{F}_{K_{2}}\right\|\leq(\|A-BK_{1}\|+\|A-BK_{2}\|)\|B\|\left\|K_{1}-K_{2}\right\|,$ for any $K_{1}$ and $K_{2}$ in $\Omega$.
Lemma 9.9 follows a direct calculation with the triangle inequality.
@@
For any $K_{1}\in\Omega$ and $K_{2}\in\Omega$, with $\xi_{\gamma,\rho}$ defined in (5.1),
(9.23) $\left\|\left(\mathcal{T}_{K_{1}}-\mathcal{T}_{K_{2}}\right)(X)\right\|$ $\leq\sum_{t=0}^{\infty}\gamma^{t}\left\|\left(\mathcal{G}_{t}^{K_{1}}-\mathcal{G}^{K_{2}}_{t}\right)(X)\right\|\leq\xi_{\gamma,\rho}\|\mathcal{F}_{K_{1}}-\mathcal{F}_{K_{2}}\|\|X\|,$
(9.24) $\sum_{t=0}^{T-1}\|(\mathcal{G}_{t}^{\prime}-\mathcal{G}_{t})(X)\|$ $\leq\frac{2-\rho^{2}-\rho^{2T}}{(1-\rho^{2})^{2}}\|\mathcal{F^{\prime}}-\mathcal{F}\|\|X\|,\quad\forall\,T\geq 1.$
Proof 9.11.
@@
(9.25) $\left\|\left(\mathcal{G}_{t+1}^{\prime}-\mathcal{G}_{t+1}\right)(X)\right\|$ $=\left\|\mathcal{F}^{\prime}\circ\mathcal{G}_{t}^{\prime}(\Sigma)-\mathcal{F}^{\prime}\circ\mathcal{G}_{t}(X)+\mathcal{F}^{\prime}\circ\mathcal{G}_{t}(X)-\mathcal{F}\circ\mathcal{G}_{t}(X)\right\|$
@@
Summing (9.25) for $t=0,1,2,\cdots$ with a discount factor $\gamma$, we have
@@
The proof of (9.23) is finished by noting $\mathcal{G}_{0}=\mathcal{F}$ and $\mathcal{G}^{\prime}_{0}=\mathcal{F}^{\prime}$.
To show (9.24), fix integer $T\geq 1.$ Summing up (9.25) for $t=0,1,\cdots,T-1$ to get
(9.26) $\sum_{t=0}^{T-1}\left\|\left(\mathcal{G}_{t+1}^{\prime}-\mathcal{G}_{t+1}\right)(X)\right\|\leq\rho^{2}\sum_{t=0}^{T-1}\left\|\left(\mathcal{G}_{t}^{\prime}-\mathcal{G}_{t}\right)(X)\right\|+\frac{1-\rho^{2T}}{1-\rho^{2}}\left\|\mathcal{F}^{\prime}-\mathcal{F}\right\|\|X\|.$
Note that the left-hand side equals
@@
Plugging the above equation into (9.26) finishes the proof of (9.24).
Proof 9.12.
(of Lemma 5.5). Denote $\mathcal{G}_{t}=\mathcal{G}^{{K_{1}}}_{t},\mathcal{G}^{\prime}_{t}=\mathcal{G}^{{K_{2}}}_{t},\mathcal{T}=\mathcal{T}_{K_{1}},\mathcal{T}^{\prime}=\mathcal{T}_{K_{2}},\mathcal{F}=\mathcal{F}_{K_{1}}$ and $\mathcal{F}^{\prime}=\mathcal{F}_{K_{2}}$ to ease the exposition. Observe that
$\|S_{K_{1},\Sigma_{1}}-S_{K_{2},\Sigma_{2}}\|\overset{(a)}{\leq}\|(\mathcal{T}-\mathcal{T}^{\prime})\mathbb{E}[x_{0}x_{0}^{\top}]\|+\|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}(\mathcal{G}_{s}^{{}^{\prime}}-\mathcal{G}_{s})(B\Sigma_{1}B^{\top}+W)\|$
$\qquad\qquad\qquad\qquad\qquad+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\|\mathcal{G}_{t-s}^{\prime}(B\Sigma_{1}B^{\top}+W)-\mathcal{G}_{t-s}^{\prime}(B\Sigma_{2}B^{\top}+W)\|$
$\overset{(b)}{\leq}\xi_{\gamma,\rho}\|\mathcal{F}-\mathcal{F}^{\prime}\|\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\sum_{t=0}^{\infty}\gamma^{t}\frac{2-\rho^{2}-\rho^{2t}}{(1-\rho^{2})^{2}}\|\mathcal{F^{\prime}}-\mathcal{F}\|\|B\Sigma_{1}B^{\top}+W\|$
$\qquad+\sum_{t=0}^{\infty}\gamma^{t}\sum_{s=1}^{t}\rho^{2(t-s)}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|,$
where $(a)$ is from (3.1) and $(b)$ follows from Lemma 9.10 and $\|\mathcal{G}_{t}(X)-\mathcal{G}_{t}(X^{\prime})\|=\|(A-BK)^{t}(X-X^{\prime})(A-BK)^{\top t}\|\leq\rho^{2t}\|X-X^{\prime}\|,$ $\forall X,X^{\prime}\in\mathbb{R}^{n\times n},\forall t\geq 0.$ Finally, applying (5.1) and Lemma 9.9 finishes the proof.
@@
The objective of this section is to provide a bound for $\|S^{(t+1)}-S^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$, as summarized in Lemma 5.6.
Note that Lemma 5.5 can be employed to derive a bound on $\|S^{(t+1)}-S^{*}\|$ in relation to $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|.$ In this section, we further establish this bound by deriving bounds for $\|K^{(t+1)}-K^{*}\|$ and $\|\Sigma^{(t+1)}-\Sigma^{*}\|$ in terms of $\|K^{(t)}-K^{*}\|$ and $\|P_{K^{(t)}}-P_{K^{*}}\|$ (cf. Lemma 9.13 and Lemma 9.15). Additionally, the perturbation analysis for $P_{K}$ in Lemma 9.17 demonstrates $\|P_{K^{(t)}}-P_{K^{*}}\|$ can be bounded by $\|{K^{(t)}}-{K^{*}}\|$, which completes the proof for Lemma 5.6.
@@
(9.28) $\|K^{\prime}-K^{*}\|$ $=$ $\|K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}-K^{*}+(R+\gamma B^{\top}P_{K}B)^{-1}E_{K^{*}}\|$
@@
(9.29) $\|E_{K}-E_{K^{*}}\|$ $\leq$ $\gamma\|B\|\|A\|\|P_{K}-P_{K^{*}}\|+\|(\gamma B^{\top}P_{K^{*}}B+R)K^{*}-(\gamma B^{\top}P_{K^{*}}B+R)K\|$
@@
Combining (9.28) and (9.29), then using $\kappa\geq\|K\|$ for any $K\in\Omega$ completes the proof.
@@
Proof 9.16.
(of Lemma 9.15). Observe from (IPO) and (2.5) that
@@
For any $K\in\Omega,$ with $c$ defined in (5.2), $\|P_{K}-P_{K^{*}}\|\leq c\|K-K^{*}\|.$
Proof 9.18.
(of Lemma 9.17). Fix $K\in\Omega.$ By (9.22) and (3.5),
@@
(9.30) $\|P_{K}-P_{K^{*}}\|\leq\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|+\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|.$
For the first term in (9.30), we can apply Lemma 9.10 and Lemma 9.9 to get
(9.31) $\left\|\mathcal{T}_{K}\left(Q+K^{\top}RK\right)-\mathcal{T}_{K^{*}}\left(Q+K^{\top}RK\right)\right\|$
@@
For the second term in (9.30), note that by Lemma 17 in [15], $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\|\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\|.$ Since $S_{K,\Sigma}\succeq\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])$, thus $\|\mathcal{T}_{K}\|\leq\frac{1}{\mu}\sigma_{max}\left(\mathcal{T}_{K}(\mathbb{E}[x_{0}x_{0}^{\top}])\right)\leq\frac{1}{\mu}\|S_{K,\Sigma}\|.$ Then we have
(9.32) $\|\mathcal{T}_{K^{*}}\|\|{K^{*}}^{\top}R{K^{*}}-K^{\top}RK\|$
@@
Plugging (9.31), (9.32), and (5.2) in (9.30) completes the proof.
@@
Proof 9.19.
(of Lemma 5.6). Let $K^{\prime}$ denote $K^{(t+1)}$ and $K$ denote $K^{(t)}$. With the assumption that $\|A-BK^{\prime}\|\leq\rho$, $\|A-BK^{*}\|\leq\rho$, we can apply Lemma 5.5 to get
(9.33) $\|$ $S_{K^{*},\Sigma^{*}}-S_{K^{\prime},\Sigma^{\prime}}\|\leq\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma^{*}-\Sigma^{\prime}\|$
$\quad+\left(\xi_{\gamma,\rho}\cdot\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\cdot\|B\Sigma^{*}B^{\top}+W\|\right)\cdot 2\rho\|B\|\left\|K^{*}-K^{\prime}\right\|.$
Apply Lemma 9.13 and Lemma 9.17 to get
$\|K^{\prime}-K^{*}\|\leq\Big($ $1+\sigma_{\min}(R)\cdot\|\gamma B^{\top}P_{K^{*}}B+R\|+c\gamma\sigma_{\min}(R)\left(\|B\|\|A\|+\|B\|^{2}\kappa\right)\Big)$
$\cdot\|K-K^{*}\|.$
Apply Lemma 9.15 and 9.17 to get $\|\Sigma^{\prime}-\Sigma^{*}\|\leq\frac{\tau\gamma\|B\|^{2}\|P_{K}-P_{K^{*}}\|}{2\sigma_{\min}(R)^{2}}\leq c\frac{\tau\gamma\|B\|^{2}\|{K}-{K^{*}}\|}{2\sigma_{\min}(R)^{2}}.$ Finally, plugging the above two inequalities into (9.33) finishes the proof.
Acknowledgement
Renyuan Xu is partially supported by the NSF CAREER Award DMS-2524465.